# Kinematics of Acceleration-Induced Excitations in Confined Quantum Fields

### v2

@@ 3 The Bogoliubov Coefficients
The two alternative representations, (24) and (27), emphasize different asymptotic regimes: the former is more convenient for large $n$, while the latter simplifies for large $m$, corresponding to high-frequency inertial or accelerated modes respectively.
@@ 3.2 Exact Solution in Terms of the Incomplete Gamma Functions
While the integral expressions above are compact, they obscure the analytic structure of the coefficients. To make this structure explicit and facilitate asymptotic analysis, it is useful to obtain a closed-form representation.
@@
In the final form, we identify the integral as the incomplete gamma function, defined by
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\times\gamma\left(-i\frac{n\pi}{\ln(1+aL)},\mp i\frac{m\pi}{aL},\mp i\frac{m\pi}{aL}\mp m\pi\right)\Bigg]$ (36)
@@
$\Omega_{n}\coloneqq\frac{n\pi a}{\ln(1+aL)}\,$,
@@
This expression makes transparent the dependence of the Bogoliubov coefficients on the ratio of mode frequencies to acceleration, showing how the confinement modifies the characteristic Unruh-type frequency scaling. Further, this form bears a striking resemblance to the Bogoliubov transformation for a scalar field propagating freely in Minkowski spacetime, when comparing inertial and uniformly accelerated observers. In the latter case, the Bogoliubov coefficients are given by
@@ 4.2 Excitation Spectrum
$P_{n}=a^{2}L^{2}\frac{4}{\pi^{4}}\sum_{\begin{subarray}{c}m=1\\ n+m=\text{odd}\end{subarray}}^{\infty}\frac{nm}{(n+m)^{6}}\,.$ (49)
@@ 5.4 Excitation Spectrum
Figure 5: Exact spectra of excitations, as calculated before, shown on a log-log plot, with the large-$n$ fit shown in dashed lines.
@@ 5.5 Excitation Spectrum in Frequency Domain
An important point now is that observers located at different positions within the accelerating box will perceive different frequencies for the same mode due to gravitational [^1] The uniform acceleration of the box is equivalent to a uniform gravitational field across the box. Hence, this time dilation can be interpreted as gravitational. Alternatively, one could just note that the Rindler metric is conformal, and the conformal factor varies with $x$. time dilation. The proper time $\tau_{O}$ for an observer at proper distance $\xi$ from the left edge of the box evolves as $d\tau_{O}=(1+a\xi)\,dt$. Consequently, the frequency of the $n^{\text{th}}$ mode as measured by the observer is
@@ 6.1 From $a_{i}$ to $a_{f}$
To start with, let’s ask: what are the Bogoliubov coefficients when a box already undergoing uniform acceleration changes its acceleration at a given instant? Suppose the left edge of the box initially accelerates at rate $a_{i}$, and this is changed to $a_{f}$ at some instant, all the while as the box maintains constant proper length. What are the resulting Bogoliubov coefficients $\alpha_{nm}$ and $\beta_{nm}$ associated with this transition?
@@
It is worth noting that for most realistic experimental setups, the dimensionless parameter $aL/c^{2}$ is extremely small. Thus, Eq.(96) and the corresponding large-$n$ excitation spectrum
$P_{n}\approx\frac{1}{10\pi^{4}}\cdot\frac{(a_{f}-a_{i})^{2}L^{2}}{n^{3}}$ (97)
are sufficient for practical analyses. (Here we are assuming that the box is in the vacuum state during its initial $a_{i}$ phase of acceleration.) Physically, the first-order $\beta$-coefficients, proportional to $a_{f}-a_{i}$, indicate that particle creation arises from changes in acceleration rather than its absolute value; if the acceleration remains constant, no new particles are produced. Consequently, the excitation spectrum in Eq. (97) scales with the square of the change in acceleration, reflecting the intensity of the perturbation imparted to the field.
@@ 7 Constraints and Challenges for Detection
We begin by noting that the excitation spectrum $P_{n}$, as given by Eq.(86), is orders of magnitude larger than what one would expect from an Unruh-type radiation effect. By an Unruh-type effect, we refer to a thermal spectrum of the form
$\Pi_{n}=\frac{1}{e^{\frac{n\pi a}{\ln(1+aL)}\cdot\frac{2\pi}{a}}-1}\,$, (103)
which corresponds to a thermal distribution with mode frequency $\Omega_{n}=n\pi a/\ln(1+aL)$ and Unruh temperature $T_{\text{Unruh}}=a/2\pi$. A comparison between $P_{n}$ and $\Pi_{n}$ for various values of acceleration is shown in Fig. 6. For the small accelerations achievable in laboratory settings, $P_{n}$ clearly dominates and thus presents a more promising candidate for probing non-inertial quantum field effects.
@@
As an example, consider a perfectly reflecting metal cavity of width $L=0.01\,\text{m}$, undergoing a change in acceleration of $100\,\text{m/s}^{2}$ (approximately 10g10g). [^2] While the electromagnetic field is not a scalar field, the mode functions for a photon is the same as for a free massless scalar field, with an added complication of polarization. We will not worry about this, and use the results above as it is for photons in a reflecting cavity. Suppose we observe photons in the $n=1000$ and $n=10000$ modes, corresponding to the infrared regime. Figure 7 shows the variation of $P_{n}/\mathcal{N}_{n}$ with temperature for these modes. To detect photons, sufficiently above the signal to (thermal) noise ratio, say near $n=10000$, the cavity must be cooled to approximately $60\,\text{K}$, while for $n=1000$, a temperature of around $5\,\text{K}$ suffices which are both well within reach of current cryogenic technology.
@@
However, if we allow ourselves to speculate beyond current technological constraints, one could imagine a future setup involving two vast $100\,\text{km}\times 100\,\text{km}$ metal sheets, separated by $1\,\text{cm}$, undergoing acceleration changes of a million $g$’s every $10^{-7}\,\text{s}$. In such a scenario, a few particles could be detected annually.
@@ 8 Discussion
The asymptotic form of the excitation spectrum, $P_{n}\sim a^{2}/n^{3}$, bears a close analogy to the Larmor formula for classical radiation, where the emitted power scales as $a^{2}$. In both classical and quantum settings, acceleration serves as a source of radiation, where the former produces electromagnetic waves, the latter generates quanta from vacuum fluctuations. This correspondence, previously noted in single-mirror setups [21], reinforces the conceptual continuity between classical and quantum radiation mechanisms [48; 49; 50; 51].

### v3

@@ 3 The Bogoliubov Coefficients
The two alternative representations, (24) and (27), emphasize different asymptotic regimes: the former is more convenient for large $n$, while the latter simplifies for large $m$, corresponding to high-frequency inertial or accelerated modes respectively. That they agree has been numerically checked as well.
@@ 3.2 Exact Solution in Terms of the Incomplete Gamma Functions
While the integral expressions of Eq.(24) and Eq.(27) obtained for the Bogoliubov coefficients are compact, they obscure the analytic structure of the coefficients. To make this structure explicit and facilitate asymptotic analysis, it is useful to obtain a closed-form representation.
@@
Here, we have substituted $w=1+az$ and $q=\frac{m\pi}{aL}w$. In the final form, we identify the integral as the incomplete gamma function, defined by
@@
$\quad\quad\quad\quad\quad\quad\quad\quad\times\gamma\left(-i\frac{n\pi}{\ln(1+aL)},\mp i\frac{m\pi}{aL},\mp i\frac{m\pi}{aL}\mp im\pi\right)\Bigg]$ (36)
The agreement of this equation with Eq.(27) has been numerically checked as well.
@@
$\Omega_{n}\coloneqq\frac{n\pi}{l}=\frac{n\pi a}{\ln(1+aL)}\,$, (37)
@@
This expression makes transparent the dependence of the Bogoliubov coefficients on the ratio of mode frequencies to acceleration, showing how the confinement modifies the characteristic Unruh-type frequency scaling. Further, this form bears a striking resemblance to the Bogoliubov transformation for a scalar field propagating freely in Minkowski spacetime, when comparing inertial and uniformly accelerated observers. In the latter case, the Bogoliubov coefficients are given by [47]
@@ 4.2 Excitation Spectrum
$P_{n}=a^{2}L^{2}\frac{4}{\pi^{4}}\sum_{\begin{subarray}{c}m=1\\ n+m=\text{odd}\end{subarray}}^{\infty}\frac{nm}{(n+m)^{6}}=\frac{4a^{2}L^{2}}{\pi^{2}}\begin{cases}\sum_{m=1}^{\infty}\frac{n(2m-1)}{n+2m-1}&\text{if }n\text{ is even}\\[6.0pt] \sum_{m=1}^{\infty}\frac{n(2m)}{n+2m}&\text{if }n\text{ is odd}\end{cases}\,.$ (50)
@@ 5.4 Excitation Spectrum
Figure 5: Exact spectra of excitations, as calculated before using Eq.(38), shown on a log-log plot, with the large-$n$ fit of Eq.(87) shown in dashed lines.
@@ 5.5 Excitation Spectrum in Frequency Domain
An important point now is that observers located at different positions within the accelerating box will perceive different frequencies for the same mode due to gravitational [^1] The uniform acceleration of the box is equivalent to a uniform gravitational field across the box. Hence, this time dilation can be interpreted as gravitational. Alternatively, one could just note that the Rindler metric is conformal, and the conformal factor varies with $x$. time dilation. The proper time $\tau_{O}$ for an observer at proper distance $\xi$ from the left edge of the box evolves as $d\tau_{O}=(1+a\xi)\,dt$. Using Eq.(37) for the frequency of the $n^{\text{th}}$ mode as on the left edge, the frequency of the $n^{\text{th}}$ mode as measured by the observer is
@@ 6.1 From $a_{i}$ to $a_{f}$
To start with, let’s ask: what are the Bogoliubov coefficients when a box already undergoing uniform acceleration changes its acceleration at a given instant? Suppose the left edge of the box initially accelerates at rate $a_{i}$, and this is changed to $a_{f}$ at some instant, all the while as the box maintains constant proper length. What are the resulting Bogoliubov coefficients $\alpha_{nm}$ and $\beta_{nm}$ associated with this transition? What would be the spectrum of excitations created in the box, if the field was in the vacuum state before the transition (in the $a_{i}$ phase of acceleration)?
@@
It is worth noting that for most realistic experimental setups, the dimensionless parameter $aL/c^{2}$ is extremely small. Thus, Eq.(97) and the corresponding spectrum of excitations for a field initially in the vacuum state for large $n$,
$P_{n}\approx\frac{1}{10\pi^{4}}\cdot\frac{(a_{f}-a_{i})^{2}L^{2}}{n^{3}}\,$, (98)
are sufficient for practical analyses. Physically, the first-order $\beta$-coefficients, proportional to $a_{f}-a_{i}$, indicate that particle creation arises from changes in acceleration rather than its absolute value; if the acceleration remains constant, no new particles are produced. Consequently, the excitation spectrum in Eq. (98) scales with the square of the change in acceleration, reflecting the intensity of the perturbation imparted to the field.
@@ 7 Constraints and Challenges for Detection
We begin by noting that the excitation spectrum $P_{n}$, as given by Eq.(87), is orders of magnitude larger than what one would expect from an Unruh-type radiation effect. By an Unruh-type effect, we refer to a thermal spectrum of the form with a temperature given by the Unruh temperature, $T_{\text{Unruh}}=a/2\pi$. Using Eq.(37) for the frequency of the modes, this thermal spectrum would look like
$\Pi_{n}=\frac{1}{e^{\frac{n\pi a}{\ln(1+aL)}\cdot\frac{2\pi}{a}}-1}\,.$ (104)
A comparison between $P_{n}$ and $\Pi_{n}$ for various values of acceleration is shown in Fig. 6. For the small accelerations achievable in laboratory settings, $aL\lesssim c^{2}$ and $P_{n}$ clearly dominates and thus presents a more promising candidate for probing non-inertial quantum field effects.
@@
As an example, consider a perfectly reflecting metal cavity of width $L=0.01\,\text{m}$, undergoing a change in acceleration of $100\,\text{m/s}^{2}$ (approximately 10g10g). [^2] While the electromagnetic field is not a scalar field, the mode functions for a photon is the same as for a free massless scalar field, with an added complication of polarization. We will not worry about this, and use the results above as it is for photons in a reflecting cavity. Suppose we observe photons in the modes $n=1000$ and n=10000n=10000333Both of these modes would lie roughly in the infrared regime. Figure 7 shows the variation of the signal to noise ratio, $P_{n}/\mathcal{N}_{n}$, with temperature for these modes. To detect photons, sufficiently clearly above the (thermal) noise, say near $n=10000$, the cavity must be cooled to approximately $60\,\text{K}$, while for $n=1000$, a temperature of around $5\,\text{K}$ is required, both of which are well within reach of current cryogenic technology.
@@ 8 Discussion
The asymptotic form of the excitation spectrum, $P_{n}\sim a^{2}/n^{3}$, bears a close analogy to the Larmor formula for classical radiation, where the emitted power scales as $a^{2}$. In both classical and quantum settings, acceleration serves as a source of radiation, where the former produces electromagnetic waves, the latter generates quanta from vacuum fluctuations. This correspondence, previously noted in single-mirror setups [21], reinforces the conceptual continuity between classical and quantum radiation mechanisms [49; 50; 51; 52].